# Carrier Localization and Spontaneous Formation of Two-Dimensional Polarization Domain in Halide Perovskites

### v1

@@ Acknowledgements.
Acknowledgments We acknowledge support from the Computational Materials Sciences Program funded by the US Department of Energy, Office of Science, Basic Energy Sciences, Materials Sciences and Engineering Division for the materials application and development of ab initio tight-binding code. We acknowledge support from the computational chemical science program within the Office of Science at DOE under grant No. DE-SC0023301 for the code development of tight-binding-LAMMPS interfaces and parallelization. Additional support for data analysis and interpretation was provided by user program of the Molecular Foundry, Lawrence Berkeley National Laboratory, supported by the Office of Science, Office of Basic Energy Sciences, of the U.S. Department of Energy under Contract No. DE-AC02-05CH11231.References

### v2

@@ Acknowledgements.
Acknowledgments We acknowledge support from the Computational Materials Sciences Program funded by the US Department of Energy, Office of Science, Basic Energy Sciences, Materials Sciences and Engineering Division for the materials application and development of ab initio tight-binding code. We acknowledge support from the computational chemical science program within the Office of Science at DOE under grant No. DE-SC0023301 for the code development of tight-binding-LAMMPS interfaces and parallelization. Additional support for data analysis and interpretation was provided by user program of the Molecular Foundry, Lawrence Berkeley National Laboratory, supported by the Office of Science, Office of Basic Energy Sciences, of the U.S. Department of Energy under Contract No. DE-AC02-05CH11231. Part of this work was performed under the auspices of the U.S. Department of Energy by Lawrence Livermore National Laboratory under Contract No. DE-AC52-07NA27344.References
@@
- Boyer-Richard et al. [2016] S. Boyer-Richard, C. Katan, B. Traoré, R. Scholz, J.-M. Jancu, and J. Even, J. Phys. Chem. Lett. 7, 3833 (2016), publisher: American Chemical Society.
- Zheng and Wang [2019] F. Zheng and L.-w. Wang, Energy Environ. Sci. 12, 1219 (2019).
- Nestoklon [2021] M. O. Nestoklon, Computational Materials Science 196, 110535 (2021).
@@
- Zhang et al. [2019] L. Zhang, D.-Y. Lin, H. Wang, R. Car, and W. E, Physical Review Materials 3, 023804 (2019).
- Zhang et al. [2018a] L. Zhang, J. Han, H. Wang, R. Car, and W. E, Physical Review Letters 120, 143001 (2018a).
- Zhang et al. [2018b] L. Zhang, J. Han, H. Wang, W. A. Saidi, R. Car, and W. E, Advances in Neural Information Processing Systems 31 , 4436 (2018b).
@@
- Xiao et al. [2020] X. Xiao, W. Li, Y. Fang, Y. Liu, Y. Shao, S. Yang, J. Zhao, X. Dai, R. Zia, and J. Huang, Nature communications 11, 2215 (2020).
- Zhang et al. [2023] B. Zhang, S. Sun, Y. Jia, J. Dai, D. T. Rathnayake, X. Huang, J. Casasent, G. Adhikari, T. A. Billy, Y. Lu, et al., Advanced Materials 35, 2208336 (2023).
- Warwick et al. [2019] A. R. Warwick, J. Íñiguez, P. D. Haynes, and N. C. Bristowe, The journal of physical chemistry letters 10, 1416 (2019).
- Jiang et al. [2025] J. Jiang, T. van der Heide, S. Thébaud, C. R. Lien-Medrano, A. Fihey, L. Pedesseau, C. Quarti, M. Zacharias, G. Volonakis, M. Kepenekian, et al., Physical Review Materials 9, 023803 (2025).
@@
- Protesescu et al. [2015] L. Protesescu, S. Yakunin, M. I. Bodnarchuk, F. Krieg, R. Caputo, C. H. Hendon, R. X. Yang, A. Walsh, and M. V. Kovalenko, Nano letters 15, 3692 (2015).
- Shimosako et al. [2024] N. Shimosako, M. Kumamoto, Y. Muroga, Z. Liu, M. Sotome, T. Kondo, H. Kunugita, and K. Ema, The Journal of Chemical Physics 161 (2024).
@@
- Bhaumik and Pandey [2024] A. Bhaumik and A. Pandey, Physical Chemistry Chemical Physics 26, 25670 (2024).
- Moualhi et al. [2022] Y. Moualhi, M. Smari, H. Rahmouni, and K. Khirouni, ACS Applied Electronic Materials 4, 4893 (2022).